# Towards Sharp Minimax Risk Bounds for Operator Learning

### v1

@@ 4.2 Error bounds for the estimator
where $1<q\leq\infty$ satisfies $1/p+1/q=1$. Now $Z_{j}$ is a sum of $N_{C}$ i.i.d. $\mathcal{N}(0,1)$ random variables. Therefore, it is subgaussian with parameters $\beta=2$ and $\kappa=1/(2N_{C})$. By a standard moment bound, we have
@@ 5.1 Lower bounds
For completeness, we establish a lower bound for $\beta>0$, as discussed in Remark 2.2. For convenience, let $\mu_{d}=\sum^{d}_{j=1}1/\lambda_{j}$ and $k=L^{2}m/(c\sigma^{2})$. Using Theorem 3.1, we can write
$\mathcal{L}_{m}(\mathcal{F};L^{p}_{\mu})$ $\lesssim\log\mathopen{}\mathclose{{\left(\frac{\sqrt{\mu_{d}}}{L}}}\right)+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\max\mathopen{}\mathclose{{\left\{\log\mathopen{}\mathclose{{\left(\frac{8}{a}}}\right),\frac{1}{2+d}\log\mathopen{}\mathclose{{\left(\frac{k}{\mu_{d}}}}\right)-\frac{d}{2+d}\log(a)}}\right\}$
$=:G(k,d)$
for all sufficiently large $d$. Now suppose that $k,d\rightarrow\infty$ with
$\frac{\log(k)-\log(\mu_{d})}{d}\sim\frac{\log(k)}{d}\text{ and }\frac{\log(k)}{d}\rightarrow\infty.$ (5.1)
@@
$G(k,d)\sim\frac{1}{2}\log(\mu_{d})+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\frac{\log(k)}{d}.$ (5.2)
We now treat the three cases $0<\beta\leq 1$ and $\beta>1$ separately. Suppose first that $0<\beta\leq 1$ and let
$d=d(k)=\mathopen{}\mathclose{{\left\lfloor\sqrt{\log(k)}}}\right\rfloor.$ (5.3)
Notice that (5.1) holds for this choice of $d$, since $\mu_{d}\leq d\exp(\alpha d^{\beta})$ and therefore $\log(\mu_{d(k)})=\mathcal{O}(\sqrt{\log(k)})=o(\log(k))$ as $k\rightarrow\infty$. With this, we have
$G(k,d(k))\asymp_{\iota,p}\sqrt{\log(k)},\quad k\rightarrow\infty.$
Writing $\log(k)=\log(m/\sigma^{2})+\log(L^{2}/c)$, we deduce that
@@
Now consider $\beta>1$. Observe that $\mu_{d}=\sum^{d}_{i=1}\exp(\alpha i^{\beta})$ satisfies
$\exp(\alpha d^{\beta})\leq\mu_{d}\leq d\exp(\alpha d^{\beta}).$
Hence, since $\beta>1$,
$G(k,d)\sim\frac{\alpha}{2}d^{\beta}+\frac{\log(k)}{d},\quad k,d\rightarrow\infty.$
@@
$d(k)=\mathopen{}\mathclose{{\left\lfloor\mathopen{}\mathclose{{\left(\log(k)}}\right)^{\frac{1}{\beta+1}}}}\right\rfloor$
@@
$G(k,d(k))\asymp_{\alpha}(\log(k))^{\frac{\beta}{\beta+1}},$
@@
$\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha}(\log(m/\sigma^{2}))^{\frac{\beta}{\beta+1}},\quad m\rightarrow\infty,$
@@
The argument is similar to the previous case. Let $\mu_{d}=\sum^{d}_{j=1}1/\lambda_{j}$, $k=L^{2}m/(c\sigma^{2})$ and use Theorem 3.1 to write
$\mathcal{L}_{m}(\mathcal{F};L^{p}_{\mu})$ $\lesssim\log\mathopen{}\mathclose{{\left(\frac{\sqrt{\mu_{d}}}{L}}}\right)+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\max\mathopen{}\mathclose{{\left\{\log\mathopen{}\mathclose{{\left(\frac{8}{a}}}\right),\frac{1}{2+d}\log\mathopen{}\mathclose{{\left(\frac{k}{\mu_{d}}}}\right)-\frac{d}{2+d}\log(a)}}\right\}$
$=:G(k,d)$
for all sufficiently large $d$. Now suppose once more that (5.1) holds, so that $G$ satisfies (5.2). Observe that $\mu_{d}=\sum^{d}_{j=1}j^{\alpha}\sim\frac{d^{\alpha+1}}{\alpha+1}$, $d\rightarrow\infty$, and therefore
$G(k,d)\sim\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\frac{\log(k)}{d}.$
@@
As noted, we currently lack tight characterizations of the minimax risk for exponentially-decaying eigenvalues with $0<\beta<1$ (see Remark 2.2) or algebraically-decaying eigenvalues for any $\alpha>1$. This stems from the factor $(\iota/(p+1))^{d}$ in 3.1. Since $\iota\leq 1$ (see Assumption 2.1), this factor is always exponentially-small in $d$. When the eigenvalues decay algebraically or exponentially with $0<\beta<1$, the decay of this term dominates in the asymptotic behaviour of the function $G(k,d)$. This is seen in the previous proofs. For exponential decay with $\beta\geq 1$ this term no longer dominates, with the consequence being a tight characterization. We believe the factor $(\iota/(p+1))^{d}$ is an artefact of the proof of Theorem 3.1. Removing it is an objective of future work.
@@ 5.2 Upper bounds
We may assume without loss of generality that $B,L\geq 1$, since this only potentially increases the size of the set $\mathcal{F}_{B,L}$. Suppose that $\lambda_{i}=\exp(-\alpha i^{\beta})$. Then
$\lambda_{1}\cdots\lambda_{d}\leq\exp(-c_{\alpha,\beta}d^{\beta+1}),$
for some constant $c_{\alpha,\beta}>0$ and
$\sum_{j>d}\lambda_{j}$ $\leq\int^{\infty}_{d}\exp(-\alpha t^{\beta})\,\mathrm{d}t.$
$=\frac{1}{\beta\alpha^{1/\beta}}\int^{\infty}_{\alpha d^{\beta}}u^{1/\beta-1}\exp(-u)\,\mathrm{d}u$
$=\frac{1}{\beta\alpha^{1/\beta}}\Gamma(1/\beta,\alpha d^{\beta}),$
@@
$\sum_{j>d}\lambda_{j}\lesssim_{\alpha,\beta}d^{1-\beta}\exp(-\alpha d^{\beta}).$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,\beta,B,L,p}d^{\frac{4}{2(p+2)}}k^{-\frac{2}{(p+2)d+4}}\exp(-c_{\alpha,\beta,p}d^{\beta})+d^{1/2-\beta/2}\exp(-c_{\alpha,\beta,p}d^{\beta})$
for some constant $c_{\alpha,\beta,p}>0$, provided
$k\geq\exp(c^{\prime}_{\alpha,\beta,p}d^{\beta+1})$
for some constant $c^{\prime}_{\alpha,\beta,p}>0$. Here we used the facts that $B,L\geq 1$ and $d\geq 1$. Set
$d=\mathopen{}\mathclose{{\left\lfloor\mathopen{}\mathclose{{\left(\frac{\log(k)}{c^{\prime}_{\alpha,\beta,p}}}}\right)^{\frac{1}{\beta+1}}}}\right\rfloor.$
@@
$\mathcal{M}_{m}(F_{B,L};L^{p}_{\mu})$ $\lesssim d\exp(-c_{\alpha,\beta,p}d^{\beta})\lesssim_{\alpha,\beta,p}(\log(m/\sigma^{2}))^{\frac{1}{\beta+1}}\exp\mathopen{}\mathclose{{\left(-c^{\prime\prime}_{\alpha,\beta,p}(\log(m/\sigma^{2}))^{\frac{\beta}{\beta+1}}}}\right),\quad m\rightarrow\infty.$
@@
$\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\gtrsim_{\alpha,\beta,p}(\log(m/\sigma^{2}))^{\frac{\beta}{\beta+1}},\quad m\to\infty,$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,\beta,B,L,p}r^{-t}+\exp\mathopen{}\mathclose{{\left(-c^{\prime\prime}_{\alpha,\beta,p}(\log(k))^{\frac{\beta}{\beta+1}}}}\right),\quad k\rightarrow\infty,$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\gtrsim_{\alpha,\beta,p}\min\mathopen{}\mathclose{{\left\{(\log(k))^{\frac{\beta}{\beta+1}},t\log(r)}}\right\}=\min\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})-\log(r)}}\right)^{\frac{\beta}{\beta+1}},t\log(r)}}\right\}.$
@@
$\log(r)\geq(\log(m/\sigma^{2}))^{\frac{\beta}{\beta+1}}.$
@@
We proceed in a similar manner. Let $\lambda_{i}=i^{-\alpha}$ and notice that
$\lambda_{1}\cdots\lambda_{d}=(d!)^{-\alpha}\leq(d/\mathrm{e})^{-\alpha d}$
@@
$\sum_{j>d}\lambda_{j}\lesssim_{\alpha}d^{1-\alpha}$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,B,L,p}dk^{-\frac{2}{(p+2)d+4}}\mathopen{}\mathclose{{\left(\frac{d}{\mathrm{e}}}}\right)^{-\frac{\alpha d}{(p+2)d+4}}+d^{\frac{1-\alpha}{2}},$
@@
$k\geq(d/\mathrm{e})^{-\alpha d/2}d^{\alpha\frac{(p+2)d+4}{2(p+2)}-\frac{4}{p+2}}.$
@@
$\log(k)\geq\frac{\alpha}{2}d+\frac{2(\alpha-2)}{p+2}\log(d).$ (5.4)
@@
$d=\mathopen{}\mathclose{{\left\lfloor\frac{4}{(p+2)+\alpha p}\frac{\log(k)}{\log(\log(k))}}}\right\rfloor$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,B,L,p}dk^{-\frac{2}{(p+2)d}}d^{-\frac{\alpha}{p+2}}+d^{\frac{1-\alpha}{2}}\lesssim_{\alpha,p}\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right)^{\frac{1-\alpha}{2}},\quad k\to\infty,$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,B,L,p}r^{-t}+\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right)^{\frac{1-\alpha}{2}},$
@@
$\mathcal{M}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\geq\min\mathopen{}\mathclose{{\left\{\frac{\alpha-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right),t\log(r)}}\right\}+\mathcal{O}_{\alpha,B,L,p}(1).$
@@
$t\log(r)\leq\frac{\alpha-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right).$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};\mathcal{L}^{p}_{\mu})\geq\frac{\alpha-1}{2}\log(\log(m/\sigma^{2}))+\mathcal{O}_{\alpha,\beta,B,L,p}(1),$

### v2

@@ 4.2 Error bounds for the estimator
where $1<q\leq\infty$ satisfies $1/p+1/q=1$. Now $Z_{j}$ is a sum of $N_{C}$ i.i.d. $\mathcal{N}(0,1)$ random variables. Therefore, it is subgaussian with parameters $\omega=2$ and $\kappa=1/(2N_{C})$. By a standard moment bound, we have
@@ 5.1 Lower bounds
For completeness, we establish a lower bound for $\omega>0$, as discussed in Remark 2.2. For convenience, let $\mu_{d}=\sum^{d}_{j=1}1/\lambda_{j}$ and $s=L^{2}m/(c\sigma^{2})$. Using Theorem 3.1, we can write
$\begin{split}\mathcal{L}_{m}(\mathcal{F};L^{p}_{\mu})&\lesssim\log\mathopen{}\mathclose{{\left(\frac{\sqrt{\mu_{d}}}{L}}}\right)+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\max\mathopen{}\mathclose{{\left\{\log\mathopen{}\mathclose{{\left(\frac{8}{a}}}\right),\frac{1}{2+d}\log\mathopen{}\mathclose{{\left(\frac{s}{\mu_{d}}}}\right)-\frac{d}{2+d}\log(a)}}\right\}\\ &=:G(s,d)\end{split}$ (5.1)
for all sufficiently large $d$. Now suppose that $s,d\rightarrow\infty$ with
$\frac{\log(s)-\log(\mu_{d})}{d}\sim\frac{\log(s)}{d}\text{ and }\frac{\log(s)}{d}\rightarrow\infty.$ (5.2)
@@
$G(s,d)\sim\frac{1}{2}\log(\mu_{d})+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\frac{\log(s)}{d}.$ (5.3)
We now treat the two cases $0<\omega\leq 1$ and $\omega>1$ separately. Suppose first that $0<\omega\leq 1$ and let
$d=d(s)=\mathopen{}\mathclose{{\left\lfloor\sqrt{\log(s)}}}\right\rfloor.$ (5.4)
Notice that (5.2) holds for this choice of $d$, since $\mu_{d}\leq d\exp(\tau d^{\omega})$ and therefore $\log(\mu_{d(s)})=\mathcal{O}(\sqrt{\log(s)})=o(\log(s))$ as $s\rightarrow\infty$. With this, we have
$G(s,d(s))\asymp_{\iota,p}\sqrt{\log(s)},\quad s\rightarrow\infty.$
Writing $\log(s)=\log(m/\sigma^{2})+\log(L^{2}/c)$, we deduce that
@@
Now consider $\omega>1$. Observe that $\mu_{d}=\sum^{d}_{i=1}\exp(\tau i^{\omega})$ satisfies
$\exp(\tau d^{\omega})\leq\mu_{d}\leq d\exp(\tau d^{\omega}).$
Hence, since $\omega>1$,
$G(s,d)\sim\frac{\tau}{2}d^{\omega}+\frac{\log(s)}{d},\quad s,d\rightarrow\infty.$
@@
$d(s)=\mathopen{}\mathclose{{\left\lfloor\mathopen{}\mathclose{{\left(\log(s)}}\right)^{\frac{1}{\omega+1}}}}\right\rfloor$
@@
$G(s,d(s))\asymp_{\tau}(\log(s))^{\frac{\omega}{\omega+1}},$
@@
$\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau}(\log(m/\sigma^{2}))^{\frac{\omega}{\omega+1}},\quad m\rightarrow\infty,$
@@
The argument is similar to the previous case. Let $\mu_{d}=\sum^{d}_{j=1}1/\lambda_{j}$, $s=L^{2}m/(c\sigma^{2})$ and use Theorem 3.1 to write
$\mathcal{L}_{m}(\mathcal{F};L^{p}_{\mu})$ $\lesssim\log\mathopen{}\mathclose{{\left(\frac{\sqrt{\mu_{d}}}{L}}}\right)+\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\max\mathopen{}\mathclose{{\left\{\log\mathopen{}\mathclose{{\left(\frac{8}{a}}}\right),\frac{1}{2+d}\log\mathopen{}\mathclose{{\left(\frac{s}{\mu_{d}}}}\right)-\frac{d}{2+d}\log(a)}}\right\}$
$=:G(s,d)$
for all sufficiently large $d$. Now suppose once more that (5.2) holds, so that $G$ satisfies (5.3). Observe that $\mu_{d}=\sum^{d}_{j=1}j^{\tau}\sim\frac{d^{\tau+1}}{\tau+1}$, $d\rightarrow\infty$, and therefore
$G(s,d)\sim\frac{d}{p}\log\mathopen{}\mathclose{{\left(\frac{p+1}{\iota}}}\right)+\frac{\log(s)}{d}.$
@@
As noted, we currently lack tight characterizations of the minimax risk for exponentially-decaying eigenvalues with $0<\omega<1$ (see Remark 2.2) or algebraically-decaying eigenvalues for any $\tau>1$. This stems from the factor $(\iota/(p+1))^{d}$ in 3.1. Since $\iota\leq 1$ (see Assumption 2.1), this factor is always exponentially-small in $d$. When the eigenvalues decay algebraically or exponentially with $0<\omega<1$, the decay of this term dominates in the asymptotic behaviour of the function $G(s,d)$. This is seen in the previous proofs. For exponential decay with $\omega\geq 1$ this term no longer dominates, with the consequence being a tight characterization. We believe the factor $(\iota/(p+1))^{d}$ is an artefact of the proof of Theorem 3.1. Removing it is an objective of future work.
@@ 5.2 Upper bounds
We may assume without loss of generality that $B,L\geq 1$, since this only potentially increases the size of the set $\mathcal{F}_{B,L}$. Suppose that $\lambda_{i}=\exp(-\tau i^{\omega})$. Then
$\lambda_{1}\cdots\lambda_{d}\leq\exp(-c_{\tau,\omega}d^{\omega+1}),$
for some constant $c_{\tau,\omega}>0$ and
$\sum_{j>d}\lambda_{j}$ $\leq\int^{\infty}_{d}\exp(-\tau t^{\omega})\,\mathrm{d}t.$
$=\frac{1}{\omega\tau^{1/\omega}}\int^{\infty}_{\tau d^{\omega}}u^{1/\omega-1}\exp(-u)\,\mathrm{d}u$
$=\frac{1}{\omega\tau^{1/\omega}}\Gamma(1/\omega,\tau d^{\omega}),$
@@
$\sum_{j>d}\lambda_{j}\lesssim_{\tau,\omega}d^{1-\omega}\exp(-\tau d^{\omega}).$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau,\omega,B,L,p}d^{\frac{4}{2(p+2)}}k^{-\frac{2}{(p+2)d+4}}\exp(-c_{\tau,\omega,p}d^{\omega})+d^{1/2-\omega/2}\exp(-c_{\tau,\omega,p}d^{\omega})$
for some constant $c_{\tau,\omega,p}>0$, provided
$k\geq\exp(c^{\prime}_{\tau,\omega,p}d^{\omega+1})$
for some constant $c^{\prime}_{\tau,\omega,p}>0$. Here we used the facts that $B,L\geq 1$ and $d\geq 1$. Set
$d=\mathopen{}\mathclose{{\left\lfloor\mathopen{}\mathclose{{\left(\frac{\log(k)}{c^{\prime}_{\tau,\omega,p}}}}\right)^{\frac{1}{\omega+1}}}}\right\rfloor.$
@@
$\mathcal{M}_{m}(F_{B,L};L^{p}_{\mu})$ $\lesssim d\exp(-c_{\tau,\omega,p}d^{\omega})\lesssim_{\tau,\omega,p}(\log(m/\sigma^{2}))^{\frac{1}{\omega+1}}\exp\mathopen{}\mathclose{{\left(-c^{\prime\prime}_{\tau,\omega,p}(\log(m/\sigma^{2}))^{\frac{\omega}{\omega+1}}}}\right),\quad m\rightarrow\infty.$
@@
$\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\gtrsim_{\tau,\omega,p}(\log(m/\sigma^{2}))^{\frac{\omega}{\omega+1}},\quad m\to\infty,$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau,\omega,B,L,p}r^{-t}+\exp\mathopen{}\mathclose{{\left(-c^{\prime\prime}_{\tau,\omega,p}(\log(k))^{\frac{\omega}{\omega+1}}}}\right),\quad k\rightarrow\infty,$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\gtrsim_{\tau,\omega,p}\min\mathopen{}\mathclose{{\left\{(\log(k))^{\frac{\omega}{\omega+1}},t\log(r)}}\right\}=\min\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})-\log(r)}}\right)^{\frac{\omega}{\omega+1}},t\log(r)}}\right\}.$
@@
$\log(r)\geq(\log(m/\sigma^{2}))^{\frac{\omega}{\omega+1}}.$
@@
We proceed in a similar manner. Let $\lambda_{i}=i^{-\tau}$ and notice that
$\lambda_{1}\cdots\lambda_{d}=(d!)^{-\tau}\leq(d/\mathrm{e})^{-\tau d}$
@@
$\sum_{j>d}\lambda_{j}\lesssim_{\tau}d^{1-\tau}$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau,B,L,p}dk^{-\frac{2}{(p+2)d+4}}\mathopen{}\mathclose{{\left(\frac{d}{\mathrm{e}}}}\right)^{-\frac{\tau d}{(p+2)d+4}}+d^{\frac{1-\tau}{2}},$
@@
$k\geq(d/\mathrm{e})^{-\tau d/2}d^{\tau\frac{(p+2)d+4}{2(p+2)}-\frac{4}{p+2}}.$
@@
$\log(k)\geq\frac{\tau}{2}d+\frac{2(\tau-2)}{p+2}\log(d).$ (5.5)
@@
$d=\mathopen{}\mathclose{{\left\lfloor\frac{4}{(p+2)+\tau p}\frac{\log(k)}{\log(\log(k))}}}\right\rfloor$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau,B,L,p}dk^{-\frac{2}{(p+2)d}}d^{-\frac{\tau}{p+2}}+d^{\frac{1-\tau}{2}}\lesssim_{\tau,p}\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right)^{\frac{1-\tau}{2}},\quad k\to\infty,$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\lesssim_{\tau,B,L,p}r^{-t}+\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right)^{\frac{1-\tau}{2}},$
@@
$\mathcal{M}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\geq\min\mathopen{}\mathclose{{\left\{\frac{\tau-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(k)}{\log(\log(k))}}}\right),t\log(r)}}\right\}+\mathcal{O}_{\tau,B,L,p}(1).$
@@
$t\log(r)\leq\frac{\tau-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right).$
@@
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};\mathcal{L}^{p}_{\mu})\geq\frac{\tau-1}{2}\log(\log(m/\sigma^{2}))+\mathcal{O}_{\tau,\omega,B,L,p}(1),$